A paced assembly line has been devised to manufacture​ calculators, as the following data​ show:

                                                                                                                       

  

LOADING...

Station

Work Element Assigned

Work Element Time​ (min)

S1

A

2.72.7

S2

​D, E

0.60.6​,

0.90.9

S3

C

3.03.0

S4

​B, F, G

0.70.7​,

0.70.7​,

0.90.9

S5

​H, I, J

0.70.7​,

0.30.3​,

1.21.2

S6

K

2.42.4

a. What is the maximum hourly output rate from this​ line?

​(Hint​:

The line can only go as fast as its slowest​ workstation.)

The maximum hourly output rate is

nothing

calculators per hour. ​(Enter your response as an​ integer.)

c. If a worker is at each station and the line operates at this maximum output​ rate, how much idle time is lost during each​ 10-hour shift?

1414

minutes per shift. ​(Enter your response as an​ integer.)
Answer #1
Station Work Element Assigned Work Element Time​ (min) Total Time per station (min)
S1 A 2.7 2.7
S2 ​D,
E
0.6​, 1.5
0.9
S3 C 3 3
S4 ​B,
F,
G
0.7​, 2.3
0.7​,
0.9
S5 ​H,
I,
J
0.7​, 2.2
0.3​,
1.2
S6 K 2.4 2.4



Answer a.

With S3 being the slowest workstation with (Total) work element time of 3min per unit. The hourly output rate of the whole line will be decided by S3

The hourly output rate from this​ line = 60 min / 3min per unit = 20 calculators per hour

Answer c

Total standard time required = S1 + S2 + S3 + S4 + S5 + S6
Cycle Time (as decided by the slowest process) = 3 minutes/ unit  
Idle time= (no. of stations)*(cycle time)-(total standard time required to assemble each unit)

  1. Total standard time required= 14.1 units (=sum of time taken as mentioned in the table at each station)
  2. Hence idle time= 6*3 - 14.1 = 3.9 minutes/unit
  3. Thus, idle time in 10 hours= for 10 hours*20 units/hour*3.9 minutes/unit= 780 minutes


    Dear Student,
    Please ask, if you have any doubts through the comment section. Do rate the answer.